OpenStudy (anonymous):

What is the solution to the system of equations? -x+2y+2z=0 -x-27-2z=0 x -z=-1

OpenStudy (anonymous):

you really should start doing these by yourself...

OpenStudy (anonymous):

i get it but i dont seem to get the last part?

OpenStudy (anonymous):

last part? elaborate.

OpenStudy (anonymous):

-1 2 2 =0 -1 -2 -2 =0 1 0 -1 =1

OpenStudy (anonymous):

wait. you have -27? in your question? is that meant to be -2y?

OpenStudy (anonymous):

yeah im sorry -2y

OpenStudy (anonymous):

your system should be -1 2 2 0 -1 -2 -2 0 1 0 -1 -1

OpenStudy (anonymous):

using your calculator or some tedious row reducing by hand we get (x,y,z)=(,0,-1,1)

OpenStudy (anonymous):

okay but how do you find or what do you do to find the answer for x,y, and z

OpenStudy (anonymous):

alrighty. this is going to take a while to explain.

OpenStudy (anonymous):

for any matrix we want to row reduce, we want to get this...

OpenStudy (anonymous):